Please with the steps:

An animal takes a breath with a volume of 5.70 L at standard temperature and pressure. If air contains a 20.9% oxygen, how many molecules of oxygen are in one breath?

.............O2 molecules


1
Expert's answer
2021-03-08T07:05:29-0500

Volume of O2 = 5.70×0.209

= 1.19 L

Moles = Volume/22.4

= 1.19/22.4

= 0.0531 mol

Number of molecules = 0.0531×6.022×1023

= [3.20×1022] molecules